OpenStudy (anonymous):

x2 - 6x + 6 = 0 Solve this quadratic equation using the quadratic formula.

OpenStudy (jackellyn):

a=1 b=-6 c=6 \[(-(-6) \pm \sqrt{(-6)^{2}-4(1)(6)}) \div (2)(1)\] \[6 \pm \sqrt{36-24} /2\] \[[6 \pm \sqrt{12}/2]\] \[3 \pm (\sqrt{12}/2)\]